Required information

[The following information applies to the questions displayed below.]

Shown below is information relating to the stockholders' equity of Reeve Corporation as of December 31, 2018:

4.5% cumulative preferred stock, $100 par $ 1,100,000
Common stock, $10 par, 380,000 shares authorized,
144,000 shares issued and outstanding
1,440,000
Additional paid-in capital: Common stock 288,000
Retained earnings (Deficit) (72,000 )
Dividends in arrears 49,500

A. How many shares of preferred stock are issued and outstanding?

B. What was the original issue price per share of common stock?

C. What is total paid-in capital?

D. Total stockholders' equity is:

E. Book value per share of common stock is: (Round to 2 decimal places.)
Answer #1

Answers

  • Requirement A

>No of preferred shares issued and outstanding = $ 1100000 / $ 100 par = 11000 shares

  • Requirement B
    >Original issue price per share of common stock = ($1440000 + 288000) / 144000 shares issued
    = 1728000 / 144000
    = $ 12 per share
  • Requirement C
    Total Paid in Capital = $ 1100000 preferred stock + $ 1440000 common stock + $ 288000 additional paid in capital
    = 1100000 + 1440000 + 288000
    = $ 2,828,000
  • Requirement ‘D’
    Total Stock holder’s equity = $ 2828000 paid in capital + Retained earnings
    = 2828000 + (-72000)
    = 2828000 – 72000
    = $ 2,756,000
  • Requirement E
    Book Value per share = ($1440000 + 288000 – 49500 dividend – 72000 retained earnings) / 144000 shares
    = 1606500 / 144000
    = $ 11.16